# Hydrostatic Newton-Cartan Membranes

### v1

@@ IV.2 Axisymmetric configurations in 3+1 flat ambient spacetime
where we have defined $\omega=-\frac{2c_{3}}{c_{0}}$. Notice that this flow corresponds to a flow with a constant spatial vorticity vector $\boldsymbol{\omega}=(0,0,\omega)$. Membranes with axisymmetric symmetry can be parametrized through the embedding mapArmas:2020 :
@@
The coordinates on the membrane will then be $(t,\rho,\phi)$. Under this embedding it is easy to find that
$\tau_{a}=\delta^{t}_{a}\,,\quad\bar{h}_{ab}$ $=\left(\sec^{2}\psi\right)\delta^{\rho}_{a}\delta^{\rho}_{b}+\rho^{2}\delta^{\phi}_{a}\delta^{\phi}_{b}\,,\quad\hat{\upsilon}^{a}=-\delta^{a}_{t}\,,\quad h^{ab}=\left(\cos^{2}\psi\right)\delta^{a}_{\rho}\delta^{b}_{\rho}+\frac{\delta^{a}_{\phi}\delta^{b}_{\phi}}{\rho^{2}}.$ (111)
@@
$u^{a}=\delta^{a}_{t}+\frac{\omega}{2}\delta^{a}_{\phi}.$ (112)
@@
$\Omega^{ab}$ $=\left(\frac{\omega}{2\rho}\cos^{2}\psi\right)\left(\delta^{a}_{\rho}\delta^{b}_{\phi}-\delta^{a}_{\phi}\delta^{b}_{\rho}\right).$ (116)
Notice that the equilibrium condition implies that temperature is constant and related to $c_{0}$ through $\frac{T_{0}}{c_{0}}$ while the chemical potential $\mu$ will be a function of $\rho$ given by
@@
The extrinsic curvature on this configuration is given by
$K_{ab}=-\left(\psi^{\prime}\sec\psi\right)\delta^{\rho}_{a}\delta^{\rho}_{b}-\left(\sin\psi\right)\delta_{a}^{\phi}\delta_{b}^{\phi}.$ (118)
Using all this information, we show in table 2 all the scalars from table 1 in terms of the embedding variables. Before finishing this section, notice that for any function $f(T,\mu)$, it will be true under the equilibrium condition that
@@
V Young-Laplace equation for axisymmetric solutions
The equation describing the shape of the membrane is given by (68) and can be considered a generalization of Young-Laplace equation. In flat ambient spacetime, it takes the form
@@ V Young-Laplace equation for axisymmetric solutions
In this section, we will focus on the immediate generalization of the case studied in Armas:2020 . This generalization will consist of keeping only the terms in the equilibrium partition function that depend only on the vorticity through their chemical potential dependence. This means keeping $\chi,a^{(1)}_{1},a^{(2)}_{1},$ and $a^{(2)}_{2}$ as arbitrary functions of the temperature and chemical potential while setting all other $a^{I}_{i}$ coefficients to zero [^8] The explicit form of the full Young-Laplace equation taking all coefficients into account is quite cumbersome and by itself it is not very instructive, so we will leave its full analysis for future work.. Therefore, the equilibrium partition function will simply look like
@@
$\left(\mathcal{T_{S}}\right)^{cd}$ $=\left[\chi+a_{1}^{(1)}K_{1}+a_{1}^{(2)}K_{1}^{2}+a_{2}^{(2)}K_{2}\right]h^{cd}$ (125)
$+\left[n+n^{(1)}_{1}K_{1}+n^{(2)}_{1}K_{1}^{2}+n^{(2)}_{2}K_{2}\right]u^{c}u^{d}$
$-2\left(a^{(1)}_{1}+2a^{(2)}_{1}K_{1}\right)h^{ac}h^{bd}K_{ab}-4a^{(2)}_{2}h^{rs}h^{ac}h^{bd}K_{ar}K_{sb},$
$\mathcal{D}^{cd}$ $=a_{1}^{(1)}h^{cd}+2a_{1}^{(2)}h^{cd}K_{1}+2a_{2}^{(2)}h^{ac}h^{bd}K_{ab}.$ (126)
@@
$\left(\mathcal{T}_{S}\right)^{ab}K_{ab}$ $=\chi K_{1}+a^{(1)}_{1}K_{1}^{2}+a_{1}^{(2)}K_{1}^{3}+a_{2}^{(2)}K_{1}K_{2}+nk_{0}+n_{1}^{(1)}K_{1}k_{0}$ (127)
$+n_{1}^{(2)}K_{1}^{2}k_{0}+n^{(2)}_{2}K_{2}k_{0}-2a_{1}^{(1)}K_{2}-4a_{1}^{(2)}K_{1}K_{2}-4a_{2}^{(2)}K_{3},$
$h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}$ $=a_{1}^{(1)}K_{2}+2a_{1}^{(2)}K_{1}K_{2}+2a_{2}^{(2)}K_{3},$ (128)
where
$k_{c}$ $=K_{cd}u^{d}\,,$ (129)
$k_{0}$ $=k_{c}u^{c}\ ,$ (130)
$K_{3}$ $=K_{ab}h^{bc}K_{cd}h^{dr}K_{rs}h^{sa}.$ (131)
Combining these two terms, we end up with
$\left(\mathcal{T_{S}}\right)^{cd}+h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}$ $=\chi K_{1}+a_{1}^{(1)}K_{1}^{2}-a_{1}^{(1)}K_{2}+\left(a_{1}^{(2)}+a_{2}^{(2)}\right)\left(K_{1}^{2}-2K_{2}\right)K_{1}$ (132)
$+nk_{0}+n_{1}^{(1)}K_{1}k_{0}+n_{1}^{(2)}K^{2}_{1}k_{0}+n_{2}^{(2)}K_{2}k_{0},$
@@
$K_{3}=\frac{3}{2}K_{1}K_{2}-\frac{1}{2}K_{1}^{3}.$ (133)
@@
$D_{c}D_{d}\mathcal{D}^{cd}$ $=h^{cd}D_{c}D_{d}a_{1}^{(1)}+2h^{cd}\left[K_{1}D_{c}D_{d}a_{1}^{(2)}+2\left(D_{c}K_{1}\right)\left(D_{d}a_{1}^{(2)}\right)+a_{1}^{(2)}D_{c}D_{d}K_{1}\right]$ (134)
@@
$=2\left(a^{(2)}_{1}+a_{2}^{(2)}\right)D^{2}K_{1}+\left(\beta^{(1)}_{1}+2\beta^{(2)}_{1}K_{1}\right)\bar{a}^{2}-\left(n^{(1)}_{1}+2n^{(2)}_{1}K_{1}\right)\left(D\cdot a\right)$
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4\left(n^{(2)}_{1}+n^{(2)}_{2}\right)a^{c}D_{c}K_{1},$
@@
where we used the Codazzi-Mainardi equation and where we defined
$\beta^{I}_{(i)}$ $=\frac{\partial n}{\partial\mu}\,,$ (135)
$D^{2}$ $=h^{cd}D_{c}D_{d}\,,$ (136)
$D\cdot a$ $=D_{c}a^{c}.$ (137)

### v2

@@ 4.2 Axisymmetric configurations in 3+1 flat ambient spacetime
where we have defined $\omega=-\frac{2c_{3}}{c_{0}}$. Notice that this corresponds to a flow with a constant spatial vorticity vector $\boldsymbol{\omega}=(0,0,\omega)$. Membranes with axisymmetric symmetry can be parametrized through the embedding map[3]:
@@
The coordinates on the membrane will then be $(t,\rho,\varphi)$. Under this embedding, it is easy to find that
$\tau_{a}$ $=\delta^{t}_{a}\,$$\quad$ $\bar{h}_{ab}$ $=\left(\sec^{2}\psi\right)\delta^{\rho}_{a}\delta^{\rho}_{b}+\rho^{2}\delta^{\varphi}_{a}\delta^{\varphi}_{b}\,$, (118)
$\hat{\upsilon}^{a}$ $=-\delta^{a}_{t}\,$$\quad$ $h^{ab}$ $=\left(\cos^{2}\psi\right)\delta^{a}_{\rho}\delta^{b}_{\rho}+\frac{\delta^{a}_{\varphi}\delta^{b}_{\varphi}}{\rho^{2}}.$
@@
$u^{a}=\delta^{a}_{t}+\frac{\omega}{2}\delta^{a}_{\varphi}.$ (119)
@@
$\Omega^{ab}$ $=\left(\frac{\omega}{2\rho}\cos^{2}\psi\right)\left(\delta^{a}_{\rho}\delta^{b}_{\varphi}-\delta^{a}_{\varphi}\delta^{b}_{\rho}\right).$ (123)
The equilibrium condition implies that temperature is constant and given by $T=\frac{T_{0}}{c_{0}}$, while the chemical potential $\mu$ will be a function of $\rho$ given by
@@
Notice that the chemical potential (124) is equal to the rotational kinetic energy per unit mass of a fluid element rotating with an angular velocity $\frac{\omega}{2}$ at a distance $\rho$ from the rotation axis. The extrinsic curvature in this configuration is given by
$K_{ab}=-\left(\psi^{\prime}\sec\psi\right)\delta^{\rho}_{a}\delta^{\rho}_{b}-\left(\sin\psi\right)\delta_{a}^{\varphi}\delta_{b}^{\varphi}.$ (125)
Using all this information, we show in table 2 all the scalars from table 1 in terms of the embedding variables. Note that the normal and tangential acceleration contributions are consistent with the centripetal acceleration of a fluid element, and note that the normal vorticity is orthogonal to the tangential acceleration. Note that for any function $f(T,\mu)$, it will be true that in equilibrium
@@
5 Young-Laplace equation for axisymmetric solutions for Canham-Helfrich-like membranes
The equation describing the shape of the membrane is given by (69) and can be considered a generalization of the Young-Laplace equation. In flat ambient spacetime, it takes the form
@@ 5 Young-Laplace equation for axisymmetric solutions for Canham-Helfrich-like membranes
In this section, we will focus on the immediate generalization of the case studied in [3]. This generalization will consist of keeping only the terms in the equilibrium partition function that depend only on the vorticity through the chemical potential; that is, we will keep only the elastic contributions. This means keeping $\chi,a^{(1)}_{1},a^{(2)}_{1}$, and $a^{(2)}_{2}$ as arbitrary functions of the temperature and chemical potential while setting all other $a^{I}_{i}$ coefficients to zero [^10] The explicit form of the full Young-Laplace equation taking all coefficients into account is quite cumbersome and by itself it is not very instructive, so we will leave its full analysis for future work.. Therefore, the equilibrium partition function will simply look like
@@
$\left(\mathcal{T_{S}}\right)^{cd}$ $=\left[\chi+a_{1}^{(1)}K+a_{1}^{(2)}K^{2}+a_{2}^{(2)}\left(K\cdot K\right)\right]h^{cd}$ (130)
$+\left[n+n^{(1)}_{1}K+n^{(2)}_{1}K^{2}+n^{(2)}_{2}\left(K\cdot K\right)\right]u^{c}u^{d}$
$-2\left(a^{(1)}_{1}+2a^{(2)}_{1}K\right)h^{ac}h^{bd}K_{ab}$
$-4a^{(2)}_{2}h^{rs}h^{ac}h^{bd}K_{ar}K_{sb},$
$\mathcal{D}^{cd}$ $=a_{1}^{(1)}h^{cd}+2a_{1}^{(2)}h^{cd}K+2a_{2}^{(2)}h^{ac}h^{bd}K_{ab}.$ (131)
@@
$\left(\mathcal{T}_{S}\right)^{ab}K_{ab}$ $=\chi K+a^{(1)}_{1}K^{2}+a_{1}^{(2)}K^{3}$ (132)
$+a_{2}^{(2)}K\left(K\cdot K\right)+n\left(a\cdot n\right)+n_{1}^{(1)}K\left(a\cdot n\right)$
$+n_{1}^{(2)}K^{2}\left(a\cdot n\right)+n^{(2)}_{2}\left(K\cdot K\right)\left(a\cdot n\right)$
$-2a_{1}^{(1)}\left(K\cdot K\right)-4a_{1}^{(2)}K\left(K\cdot K\right)-4a_{2}^{(2)}K_{3},$
$h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}$ $=a_{1}^{(1)}\left(K\cdot K\right)+2a_{1}^{(2)}K\left(K\cdot K\right)$ (133)
$+2a_{2}^{(2)}K_{ab}K_{cd}K_{rs}h^{bc}h^{dr}h^{sa},$
where we have used that $K_{cd}u^{c}u^{d}=\left(a\cdot n\right)$. Combining (132) and (133), we end up with
$\left(\mathcal{T_{S}}\right)^{cd}+h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}$ $=\chi K+a_{1}^{(1)}K^{2}$ (134)
$-a_{1}^{(1)}\left(K\cdot K\right)$
$+\left[a_{1}^{(2)}+a_{2}^{(2)}\right]\left[K^{2}-2\left(K\cdot K\right)\right]K$
$+n\left(a\cdot n\right)+n_{1}^{(1)}K\left(a\cdot n\right)$
$+n_{1}^{(2)}K^{2}\left(a\cdot n\right)$
$+n_{2}^{(2)}\left(K\cdot K\right)\left(a\cdot n\right),$
@@
$K_{ab}h^{bc}K_{cd}h^{dr}K_{rs}h^{sa}=\frac{3}{2}K\left(K\cdot K\right)-\frac{1}{2}K^{3}.$ (135)
@@
$D_{c}D_{d}\mathcal{D}^{cd}$ $=h^{cd}D_{c}D_{d}a_{1}^{(1)}+2h^{cd}\left[KD_{c}D_{d}a_{1}^{(2)}+2\left(D_{c}K\right)\left(D_{d}a_{1}^{(2)}\right)\right$. (136)
$\left.+a_{1}^{(2)}D_{c}D_{d}K\right]$
@@
$=2\left(a^{(2)}_{1}+a_{2}^{(2)}\right)D^{2}K+\left(\beta^{(1)}_{1}+2\beta^{(2)}_{1}K\right)\bar{a}^{2}$
$-\left(n^{(1)}_{1}+2n^{(2)}_{1}K\right)\left(D\cdot a\right)$
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4\left(n^{(2)}_{1}+n^{(2)}_{2}\right)a^{c}D_{c}K,$
@@
where we used the shorthand notation $D^{2}=h^{cd}D_{c}D_{d}$, $D\cdot a=D_{c}a^{c}$, the Codazzi-Mainardi equation, and where we defined
$\beta^{(1)}_{i}$ $=\frac{\partial n^{(1)}_{i}}{\partial\mu}\,,\qquad\beta^{(2)}_{i}=\frac{\partial n^{(2)}_{i}}{\partial\mu}.$ (137)